Prove Lorentz conoriance of the first equation of the electromagnetic field.
divB=0, and curlE= -db/dt
1
Expert's answer
2020-05-05T18:46:27-0400

We have:


"\text{div}\textbf{B}=0,\\
\text{curl}\textbf{ E}=-\frac{\partial \textbf{B}}{\partial t}"

The first equation does not have time derivatives, we can satisfy this constraint by introducing a vector potential A which will comply with the condition


"\textbf{B}=\text{curl}\textbf{ A}."

When we set


"\text{curl}\textbf{ E}=-\frac{\partial \textbf{A}}{\partial t},"

it implies a basic law of electromagnetism, Faraday's law:


"\text{curl}\textbf{ E}=-\frac{\partial \textbf{B}}{\partial t}."

Reference

Michael Stone, Paul Goldbart (2009). Mathematics for Physics: A Guided Tour for Graduate Students. Publisher: Science.